If a baseball player has a batting average of 0.265, what is the probability that the player will get the following number of hits in the next four times at bat?
A- exactly 2 hits=
B- at least 2 hits=

Answer by Boreal(15235)

exactly is 4C2 (6 ways it can happen)*0.265^2*0.735^2=0.2276
also 2ndVARS binompdf(4,0.265,2)ENTER
-
at least 2 means 2,3,4 or 1-0,1
0 is .735^4=0.2918
1 is 4*0.735^3*0.265=0.4209
1- those is 0.2873.
Calculator 1-binompdf(4,0.265,1)ENTER